# Whisper D-SGD: Correlated Noise Across Agents for Differentially Private Decentralized Learning

### v1

@@ IV-B Privacy Analysis
We show that our framework ensures agent-level $(\varepsilon,\delta)$-DP for arbitrary covariance $\mathbf{R}\succ 0$. The following theorem bounds $\varepsilon$ in terms of the number of iterations $T$, clipping threshold $C$, and the inverse covariance $\mathbf{R}^{-1}$.
@@
Our covariance-based framework satisfies ($\varepsilon,\delta$)-DP with:
@@
We adapt existing Rényi DP (RDP) arguments from [12, 18] to arbitrary $\mathbf{R}\succ 0$. At each iteration, our framework satisfies $(\alpha,\alpha\varepsilon_{\text{step}})$-RDP with $\varepsilon_{\text{step}}=2C^{2}\max_{i\in\mathcal{N}}[\mathbf{R}^{-1}]_{ii}$ [18, Thm. 6]. Since RDP composes additively over $T$ iterations [12, Prop. 1], the final iterate $\tilde{\mathbf{x}}^{(T)}$ satisfies ($\alpha,T\alpha\varepsilon_{\text{step}}$)-RDP. Converting RDP to $(\varepsilon,\delta)$-DP introduces an additional term $\ln(1/\delta)/(\alpha-1)$ [12, Prop. 3]. Optimizing the expression over $\alpha>1$ concludes the proof. ∎
Theorem 2 directly relates the diagonal elements of $\mathbf{R}^{-1}$, specifically $\max_{i\in\mathcal{V}}[\mathbf{R}^{-1}]_{ii}$, to the per-agent privacy budget $\varepsilon$. This result immediately recovers well-known LDP bounds (e.g., [12, Prop. 7]) when $\mathbf{R}=\sigma_{\text{ldp}}^{2}\mathbf{I}_{n}$: larger $[\mathbf{R}]_{ii}$ (i.e., larger variance of privacy noise) implies smaller $[\mathbf{R}^{-1}]_{ii}$, thus smaller $\varepsilon$ (that is, strengthened $(\varepsilon,\delta)$-DP guarantees).
Invertibility ($\mathbf{R}\succ 0$) is a necessary condition for ensuring agent-level privacy under correlated noise: if $\mathbf{R}$ were singular, linear combinations of noise could cancel completely among agents, directly exposing their local models.
Both the number of iterations $T$ and the clipping threshold $C$ appear as multiplicative factors in Eq. (8), increasing $\varepsilon$. Intuitively, more rounds of communication or higher gradient sensitivity (larger $C$) amplify privacy loss, requiring more noise to protect local models.
V The “Whisper D-SGD” Algorithm
@@ V-A Privacy-Utility Trade-Off
(G1) Invertibility. A necessary condition for LDP-level privacy is $\mathbf{R}\succ 0$ (Theorem 2). For this reason, we partition $\mathbf{R}$ into an independent noise component ($\sigma_{\text{mix}}^{2}\mathbf{I}_{n}$), that ensures invertibility, and an arbitrary correlated component ($\mathbf{R}_{\text{cor}}$):
@@
Problem (10) can be written as a convex semidefinite program (SDP). The objective is linear in $\mathbf{R}_{\text{mix}}$, and the mapping $\mathbf{R}_{\text{mix}}\mapsto\mathbf{R}_{\text{mix}}^{-1}$ is matrix-convex on $\mathbf{R}_{\text{mix}}\succ 0$ [23, §4.6.2]. By applying Schur-complement constraints, each bound on $[\mathbf{R}_{\text{mix}}^{-1}]_{ii}$ becomes a linear matrix inequality, allowing solution by standard SDP packages (e.g., MOSEK [24]). Solving (10) yields $\sigma_{\text{mix}}^{2,\star}$ and $\mathbf{R}_{\text{cor}}^{\star}$. We then compute $\mathbf{R}_{\text{mix}}^{\star}=\sigma_{\text{mix}}^{2,\star}\mathbf{I}_{n}+\mathbf{R}_{\text{cor}}^{\star}$.
@@
Let $\varepsilon\leq\log(1/\delta)$. Problem (10) maximizes the utility of our covariance-based framework under $(\varepsilon,\delta)$-DP.
@@
First, the objective in (10a) minimizes the privacy-noise variance in Theorem 1, Eq. (7). Second, applying Eq. (10b) to the privacy bound in Theorem 2, Eq. (8) shows that our framework is ($\varepsilon^{*},\delta$)-DP, where $\varepsilon^{*}\leq\frac{\varepsilon^{2}}{8\log(1/\delta)}+\frac{\varepsilon}{\sqrt{2}}\leq\varepsilon$. ∎
V-B Whisper D-SGD
Algorithm 1 Whisper D-SGD
@@ V-B Whisper D-SGD
We now introduce Whisper D-SGD (Algorithm 1), a novel differentially private D-SGD algorithm which combines our covariance-based framework and our theoretical guidelines to achieve near-CDP utility and LDP-level privacy.
Conceptual overview
In Whisper D-SGD, each agent contributes a “whisper” of noise (with variance $\sigma_{\text{mix}}^{2,\star}$) sufficient to protect its local model, thus to ensure LDP-level privacy. By design, correlated noise (with covariance $\mathbf{R}_{\text{cor}}^{\star}$) cancels during mixing, so that the final models reap many of the performance benefits typically associated with CDP-level approaches.
Implementation details
(i) Initialize $(\sigma_{\text{mix}}^{2},\mathbf{R}_{\text{cor}})$ so that $\mathbf{R}_{\text{mix}}$ is invertible (line 2). (ii) Optimize Problem (10) via standard SDP solvers (line 3). (iii) Each agent shares the optimal $\mathbf{R}_{\text{mix}}^{\star}$ and a global random seed $s$ (line 4). (iv) Execute the standard differentially-private D-SGD algorithm (lines 510), except agents add privacy noise $\mathbf{v}^{(t)}=\mathbf{R}_{\text{mix}}^{\star 1/2}\mathbf{s}^{(t)}$, where $\mathbf{s}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ is locally generated from seed $s$ (line 7).
Practical considerations
Once agents share $\mathbf{R}_{\text{mix}}^{\star}$ and seed $s$, Whisper D-SGD runs fully decentralized, incurring no extra communication overhead than baseline D-SGD (lines 510). However, computing $\mathbf{R}_{\text{mix}}^{\star}$ requires either a centralized or a distributed procedure (lines 24). In the centralized approach, a single node with knowledge of $\mathbf{W}$ and parameters $(\varepsilon,\delta,T,C)$ solves (10) and broadcasts $(\mathbf{R}_{\text{mix}}^{\star},s)$ to all agents. Alternatively, a distributed setup can achieve the same outcome by running an initial consensus phase on $\mathbf{W}$ and $s$, then locally solving (10). This approach removes reliance on any single node but incurs a one-time communication overhead.
Further extensions
While Whisper D-SGD outperforms pairwise and fully independent approaches against external eavesdroppers (Sections V-CVI), it currently requires sharing $(\mathbf{R}_{\text{mix}}^{\star},s)$ among all agents, offering no defense against honest-but-curious internal participants. Future work may distribute different seeds among subgroups of agents or impose structural constraints on $\mathbf{R}_{\text{cor}}$ to relax global trust. Another direction is the joint optimization of $\mathbf{W}$ and $\mathbf{R}$–rather than fixing $\mathbf{W}$–adapting the mixing weights to further amplify noise cancellation and improve convergence speed. Our framework can also be extended to directed graphs (e.g., by defining asymmetric Laplacians) and to other D-SGD variants (e.g., the push-sum algorithm [7]). Finally, if an agent departs mid-training, $\mathbf{R}_{\text{mix}}^{\star}$ must be currently recomputed. Resilient algorithms to update $\mathbf{R}_{\text{mix}}$ on the fly remain an open challenge.
@@ V-C Comparison with Related Works
(a) $p=0.5$.
(b) $\varepsilon=10$.
Figure 1: Comparison of variances $\sigma_{\text{mix}}^{2}$, $\sigma_{\text{pair}}^{2}$, $\sigma_{\text{ldp}}^{2}$, $\sigma_{\text{cdp}}^{2}$ versus (a) privacy budget $\varepsilon$ at $p{=}0.5$ and (b) connectivity $p$ at $\varepsilon{=}10$.
The performance of differentially private D-SGD hinges on the independent portion of each agent’s noise, i.e., the variance that remains after any correlation-based cancellation.
To evaluate our correlated-noise design, we compare $\sigma_{\text{mix}}^{2}$ (the effective independent variance in Whisper D-SGD) against the following state-of-the-art baselines: (i) pairwise correlation, by replacing $\mathbf{R}_{\text{mix}}$ in (10) with $\mathbf{R}_{\text{pair}}=\sigma_{\text{pair}}^{2}\mathbf{I}_{n}+\sigma_{\text{cor}}^{2}\mathbf{L}$, (ii) LDP, with $\mathbf{R}_{\text{ldp}}=\sigma_{\text{ldp}}^{2}\mathbf{I}_{n}$, and (iii) CDP: $\sigma_{\text{cdp}}^{2}=\sigma_{\text{ldp}}^{2}/{n}$.
Figure 1 reports solutions of Problem (10), comparing $\sigma_{\text{mix}}^{2}$, $\sigma_{\text{pair}}^{2}$, $\sigma_{\text{ldp}}^{2}$, and $\sigma_{\text{cdp}}^{2}$ under varying privacy budgets $\varepsilon$ (Fig. 1(a)) and network connectivity $p$ in an Erdős–Rényi graph (Fig. 1(b)).
The fully independent LDP noise imposes the highest variance $\sigma_{\text{ldp}}^{2}$, as it forgoes any correlation-based cancellation (it serves as a practical upper bound). Conversely, $\sigma_{\text{cdp}}^{2}$ acts as a theoretical lower bound, representing the ideal CDP-level variance. As $\varepsilon$ and $p$ increase (weaker privacy requirements or denser connectivity), correlation-based schemes ($\sigma_{\text{pair}}^{2}$ and $\sigma_{\text{mix}}^{2}$) become more effective, with Whisper D-SGD consistently achieving $\sigma_{\text{mix}}^{2}\ll\sigma_{\text{pair}}^{2}$. This gap is particularly large for sparser networks (lower $p$), since $\mathbf{R}_{\text{mix}}$ leverages the entire mixing matrix $\mathbf{W}$, rather than just the graph Laplacian $\mathbf{L}$. Overall, Whisper D-SGD brings LDP-level privacy (through $\sigma_{\text{mix}}^{2}$) much closer to the ideal CDP regime ($\sigma_{\text{cdp}}^{2}$).
VI Experimental Evaluation
@@
We evaluate Whisper D-SGD on synthetic quadratic objectives and real-world logistic regression, under varying privacy budgets $\varepsilon$ and network connectivity $p$. Our code is available at https://github.com/arodio/WhisperDSGD.
@@ VI-A1 Network and DP parameters
We simulate $n=20$ agents on an Erdős–Rényi graph with varying connectivity $p\in\{0.2,0.4,0.6,0.8,1.0\}$. We adopt Metropolis–Hastings weights $w_{ij}=\tfrac{1}{|\mathcal{N}_{i}|+1}$ if $j\in\mathcal{N}_{i}$ [19]. Following standard benchmarks [18], we set DP parameters $\varepsilon\in\{3,5,7,10,15,20,25,30,40\}$, $\delta=10^{-5}$, and clipping threshold $C=0.1$.

### v2

@@ IV-B Privacy Analysis
We provide $(\varepsilon,\delta)$-LDP guarantees for our framework with arbitrary $\mathbf{R}\succ 0$. Theorem 2 bounds $\varepsilon$ in terms of number of iterations $T$, clipping threshold $C$, and inverse covariance $\mathbf{R}^{-1}$.
@@
Our covariance-based framework is ($\varepsilon,\delta$)-LDP:
@@
We adapt existing Rényi DP (RDP) arguments from [12, 18] to arbitrary $\mathbf{R}\succ 0$. At each iteration, our framework satisfies $(\alpha,\alpha\varepsilon_{\text{step}})$-RDP with $\varepsilon_{\text{step}}=2C^{2}\max_{i\in\mathcal{V}}[\mathbf{R}^{-1}]_{ii}$ [18, Thm. 6]. Since RDP composes additively over $T$ iterations [12, Prop. 1], the final iterate $\tilde{\mathbf{x}}^{(T)}$ satisfies ($\alpha,T\alpha\varepsilon_{\text{step}}$)-RDP. Converting RDP to $(\varepsilon,\delta)$-DP introduces an additional term $\ln(1/\delta)/(\alpha-1)$ [12, Prop. 3]. Optimizing the expression over $\alpha>1$ concludes the proof. ∎
Theorem 2 directly relates the diagonal elements of $\mathbf{R}^{-1}$, specifically $\max_{i\in\mathcal{V}}[\mathbf{R}^{-1}]_{ii}$, to the privacy budget $\varepsilon$. This result immediately recovers well-known LDP bounds (e.g., [12, Prop. 7]) when $\mathbf{R}=\sigma_{\text{ldp}}^{2}\mathbf{I}_{n}$: larger $[\mathbf{R}]_{ii}$ (i.e., larger variance of privacy noise) implies smaller $[\mathbf{R}^{-1}]_{ii}$, thus smaller $\varepsilon$ (that is, strengthened $(\varepsilon,\delta)$-LDP guarantees).
Invertibility ($\mathbf{R}\succ 0$) is a necessary condition for LDP guarantees under correlated noise: if $\mathbf{R}$ were singular, linear combinations of noise could cancel completely across agents, directly exposing their local models after the mixing step.
V The “CorN-DSGD” Algorithm
(a) $n=20$, $p=0.5$.
(b) $n=20$, $\varepsilon=10$.
(c) $p=0.5$, $\varepsilon=10$.
Figure 1: $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top})$, $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{pair}}\mathbf{W}^{\top})$, $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{ldp}}\mathbf{W}^{\top})$ vs. varying: (a) privacy budget $\varepsilon$ at $p=0.5$, (b) connectivity $p$ at $\varepsilon=10$, and (c) number of agents $n$ at $p=0.5$, $\varepsilon=10$.
(d) $n=20$, $p=0.5$, $\varepsilon=10$.
Figure 2: $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top})$ vs. number of independent seeds.
@@ V-A Privacy-Utility Trade-Off
(G1) Invertibility. A necessary condition for $(\varepsilon,\delta)$-LDP is $\mathbf{R}\succ 0$ (Theorem 2). For this reason, we partition $\mathbf{R}$ into an independent noise component ($\sigma_{\text{mix}}^{2}\mathbf{I}_{n}$), that ensures invertibility, and an arbitrary correlated component ($\mathbf{R}_{\text{cor}}$):
@@
The covariance $\mathbf{R}_{\text{mix}}$ in (8) generalizes prior approaches as special cases: (i) LDP: $\mathbf{R}_{\text{cor}}=\mathbf{0}$; (ii) DECOR: $\mathbf{R}_{\text{cor}}=\sigma_{\text{cor}}^{2}\mathbf{L}$. We remark $\sigma_{\text{mix}}^{2}>0$; if not, $\mathbf{R}_{\text{mix}}$ would be singular, and $\mathbf{v}\sim\mathcal{N}(\mathbf{0},\mathbf{R}_{\text{mix}})$ could make $\mathbf{W}\mathbf{v}=\mathbf{0}$, i.e., no privacy noise.
@@
Problem (9) can be formulated as a convex semidefinite program (SDP). The objective is linear in $\mathbf{R}_{\text{mix}}$, and the mapping $\mathbf{R}_{\text{mix}}\mapsto\mathbf{R}_{\text{mix}}^{-1}$ is matrix-convex on $\mathbf{R}_{\text{mix}}\succ 0$ [20, §4.6.2]. By applying Schur-complement constraints, each bound on $[\mathbf{R}_{\text{mix}}^{-1}]_{ii}$ becomes a linear matrix inequality. Thus, Problem (9) can be solved efficiently by standard SDP solvers (e.g., MOSEK [21]). Also, our SDP formulation is scalable, as we verify numerically with networks of up to 100 agents (Fig. 1(c)). Solving (9) yields the optimal solutions $\sigma_{\text{mix}}^{2,\star}$ and $\mathbf{R}_{\text{cor}}^{\star}$, from which we compute $\mathbf{R}_{\text{mix}}^{\star}=\sigma_{\text{mix}}^{2,\star}\mathbf{I}_{n}+\mathbf{R}_{\text{cor}}^{\star}$.
@@
Problem (9) attains the largest utility among all covariances satisfying (8); every such solution is $(\varepsilon,\delta)$-LDP.
@@
First, the objective in (9a) minimizes the privacy-noise variance in Theorem 1, Eq. (6). Second, applying Eq. (9b) to the privacy bound in Theorem 2, Eq. (7) shows that our framework is ($\varepsilon^{*},\delta$)-LDP, where $\varepsilon^{*}\leq\frac{\varepsilon^{2}}{8\log(1/\delta)}+\frac{\varepsilon}{\sqrt{2}}\leq\varepsilon$. ∎
V-B Correlated Noise DSGD (CorN-DSGD)
Algorithm 1 CorN-DSGD
@@ V-B Correlated Noise DSGD (CorN-DSGD)
We now present CorN-DSGD (Algorithm 1), a novel differentially private DSGD algorithm that integrates our covariance-based framework with the theoretical guidelines.
CorN-DSGD: (i) initializes $(\sigma_{\text{mix}}^{2},\mathbf{R}_{\text{cor}})$ so that $\mathbf{R}_{\text{mix}}$ is invertible (line 2); (ii) optimizes Problem (9) via standard SDP solvers (line 3); (iii) shares the optimal $\mathbf{R}_{\text{mix}}^{\star}$ and a global random seed $s$ among all agents (line 4); (iv) executes the standard differentially-private DSGD algorithm (lines 510), except agents add privacy noise $\mathbf{v}^{(t)}=\mathbf{R}_{\text{mix}}^{\star 1/2}\mathbf{s}^{(t)}$, where $\mathbf{s}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ is locally generated from seed $s$ (line 7).
Once agents share $\mathbf{R}_{\text{mix}}^{\star}$ and seed $s$, CorN-DSGD runs fully decentralized, incurring no extra communication overhead than baseline DSGD (lines 510). However, computing $\mathbf{R}_{\text{mix}}^{\star}$ requires either a centralized or a distributed procedure (lines 24). In the centralized approach, a single node with knowledge of $\mathbf{W}$ and parameters $(\varepsilon,\delta,T,C)$ solves (9) and broadcasts $(\mathbf{R}_{\text{mix}}^{\star},s)$ to all agents. Alternatively, a distributed setup involves an initial consensus phase on $\mathbf{W}$ and $s$, then each agent locally solves (9). This approach removes reliance on any single node but incurs an initial communication overhead.
@@ V-C Comparison with Related Works
(a) MNIST, $p=0.5$.
(b) MNIST, $\varepsilon=10$.
Figure 5: Image Classification. Test accuracy versus (a) privacy budget $\varepsilon$ at $p=0.5$ and (b) connectivity $p$ at $\varepsilon=10$.
We compare $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top})$ (the effective noise variance in CorN-DSGD) against the following state-of-the-art baselines: (i) pairwise correlation, by replacing $\mathbf{R}_{\text{mix}}$ in (9) with $\mathbf{R}_{\text{pair}}=\sigma_{\text{pair}}^{2}\mathbf{I}_{n}+\sigma_{\text{cor}}^{2}\mathbf{L}$, and (ii) LDP, with $\mathbf{R}_{\text{ldp}}=\sigma_{\text{ldp}}^{2}\mathbf{I}_{n}$.
Figure 2 reports solutions of Problem (9) for $\mathbf{R}_{\text{mix}}$, $\mathbf{R}_{\text{pair}}$, and $\mathbf{R}_{\text{ldp}}$ in an Erdős–Rényi graph under varying privacy budgets $\varepsilon$ (Fig. 1(a)), network connectivity $p$ (Fig. 1(b)), and number of agents $n$ (Fig. 1(c)). The fully independent LDP noise imposes the highest variance $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{ldp}}\mathbf{W}^{\top})$, as it forgoes any correlation-based cancellation. As $\varepsilon$ and $p$ increase (weaker privacy requirements or denser connectivity), correlation-based schemes ($\mathbf{R}_{\text{pair}}$ and $\mathbf{R}_{\text{mix}}$) become more effective, and $\mathbf{R}_{\text{mix}}$ consistently achieves lower variance than $\mathbf{R}_{\text{pair}}$. This gap is particularly large for sparser networks (lower $p$), since $\mathbf{R}_{\text{mix}}$ leverages broader covariance structures (through $\mathbf{R}_{\text{cor}}$), rather than just the graph Laplacian $\mathbf{L}$.
V-D Honest-But-Curious (HBC) Setting
We extend Problem (9) to account for honest-but-curious (HBC) agents. Let $\mathcal{S}=\{s_{k}\}_{k=1}^{m}$ be $m$ independent seeds. Seed $s_{k}$ is known to the (possibly overlapping) set of agents $\mathcal{H}_{k}\subseteq\mathcal{V}$. The covariance is decomposed as: $\mathbf{R}_{\text{cor}}=\sum_{k=1}^{m}\mathbf{R}_{\text{cor}}^{(k)}$, with $\mathrm{supp}(\mathbf{R}_{\text{cor}}^{(k)})\subseteq\mathcal{H}_{k}\times\mathcal{H}_{k}$. Thus the pair $(\mathbf{R}_{\text{cor}}^{(k)},s_{k})$ is hidden from every agent outside $\mathcal{H}_{k}$.
For any adversarial coalition $I\subseteq\mathcal{V}$, define the set of unknown seeds $\mathcal{U}(I)\coloneqq\{k:\mathcal{H}_{k}\cap I=\emptyset\}$. The covariance unknown to $I$ is $\mathbf{R}_{\text{mix}}^{(I)}\coloneqq\sigma_{\text{mix}}^{2}\mathbf{I}_{n}+\sum_{k\in\mathcal{U}(I)}\mathbf{R}_{\text{cor}}^{(k)}$. We protect against coalitions of size at most $q$: $\mathcal{C}_{q}\coloneqq\{I\subseteq\mathcal{V}:|I|\leq q\}$.
In this context, the utility analysis in Theorem 1 remains unchanged, while Theorem 2 requires only minor adaptation, as each coalition $I\in\mathcal{C}_{q}$ imposes its own privacy constraint:
$\textstyle\underset{\sigma_{\text{mix}}^{2}>0,\{\mathbf{R}_{\text{cor}}^{(k)}\}}{\text{minimize}}\quad$ $\mathrm{Tr}\bigl{(}\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top}\bigr{)}$ (10a)
subject to $\textstyle\mathbf{R}_{\text{mix}}=\sigma_{\text{mix}}^{2}\mathbf{I}_{n}+\sum_{k=1}^{m}\mathbf{R}_{\text{cor}}^{(k)},$ (10b)
$\textstyle\mathbf{R}_{\text{cor}}^{(k)}\succeq\mathbf{0},~{}\mathrm{supp}(\mathbf{R}_{\text{cor}}^{(k)})\subseteq\mathcal{H}_{k}\times\mathcal{H}_{k},$ (10c)
$\textstyle[(\mathbf{R}_{\text{mix}}^{(I)})^{-1}]_{ii}\leq\frac{\varepsilon^{2}}{16C^{2}T\log(1/\delta)},$ (10d)
$\forall I\in\mathcal{C}_{q},i\not\in I.$
Problem (10) is solvable in polynomial time. Special cases are: (i) external eavesdropper ($q=0$, $m=1$); (ii) HBC without collusion ($q=1$); (iii) DECOR ($q\geq 1$, $m=|\mathcal{E}|$); (iv) LDP ($q=n-1$, $m=n$). Fig. 2 shows the value of $\mathrm{Tr}(\mathbf{W}\mathbf{R}_{\text{mix}}\mathbf{W}^{\top})$ for different numbers of independent seeds ($m\in\{1,\dots,n\}$) and disjoint groups of agents ($\bigcap_{k}\mathcal{H}_{k}=\emptyset$).
VI Experimental Evaluation
@@ VI-A1 Network and DP parameters
We simulate $n=20$ agents on an Erdős–Rényi graph with varying connectivity $p\in\{0.2,0.4,0.6,0.8,1.0\}$. We use Metropolis–Hastings weights: $w_{ij}=\tfrac{1}{|\mathcal{N}_{i}|+1}$ if $j\in\mathcal{N}_{i}$ [19]. For comparison with [18], we focus on external eavesdroppers observing all agents’ communications ($q=0$, $m=1$), and we set DP parameters $\varepsilon\in\{3,5,7,10,15,20,25,30,40\}$, $\delta=10^{-5}$, $C=0.1$.